1. An object is 70 µm long and 47.66 µm wide. How long and wide is the object in km? (4)

A) 70 000 km long and 47622 km wide.         

B) 7.0 x 10-8 km long and 4.8 km wide 

C) 7.0 x 10-8 km long and 4.766 x 10-8 km wide

D) 0.7 x 10-8 km long and 4.766 x 10-7 km wide 


2. Which one of the following is the longest length? (4)

A) 100 decimeters C) 105 millimeters E) 10-7 nanometers

B) 10-2 centimeters D) 101 meters




1
Expert's answer
2021-06-07T09:34:10-0400

1. Converting to km, have "70\mu m = 70\times 10^{-6}m = 70\times 10^{-9}km" and "47.66\mu m = 47.66\times 10^{-6}m = 47.66\times 10^{-9}km = 4.766\times 10^{-8}km"

Thus, the answer is C.


2. D